Determine the minimum concentration of the precipitating agent on the right to cause precipitation of the cation from the solution on the left. 4.0×10−2 M Ba(NO3)2;NaF Express your answer using two significant figures 8.5×10−2 M CaI2;K2SO4 Express your answer using two significant figures 1.7×10−3 M AgNO3;RbCl Express your answer using two significant figures.
Answer #1

Solution-

1) Ksp = [Ba 2+][F-]^2 ; where Ksp of BaF2 = 2.45 * 10^-5
So, 2.45*10^-5 = (4.0*10^-2) * [F-]^2
sqrt(0.0006125) = [F-]
=0.024


2) Ksp = [Ca 2+][SO4 2-] ; where Ksp of CaSO4 = 7.10 * 10^-5
So, 7.10*10^-5 = (8.5*10^-2) * [SO4 2-]
(7.10*10^-5)/(8.5*10^-2) = [SO4 2-]

= 8.35 * 10^-4


3) Ksp = [Ag+][Cl-] ; where Ksp of AgCl = 1.77 * 10^-10
So, 1.77*10^-10 = (1.7*10^-3) * [Cl-]
(1.77*10^-10)/(1.77*10^-3) = [Cl-]

= 1*10^-7